[Halld-jlab] Hall D Staff Meeting Notes, May 5, 2014

Mark Ito marki at jlab.org
Tue May 6 20:00:18 EDT 2014


Friends,

Please find the notes below and at

https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_June_5,_2014 
.

   -- Mark
____________________________________________________

Private:Hall D Staff Meeting, June 5, 2014

Announcements, Eugene

      * Beam: goal for today is to get beam to 5C. No trace of progress on
        web pages.
           + open for controlled access today, try again for beam tomorrow
             afternoon
           + 0.6 GeV per linac to %C dumplet done
           + machine now at 900 MeV per linac
      * progress for last month good
           + 5 months to go
           + not overspending recently, short of manpower
           + spending to reduce manpower demand is good

NSAC Long-Range Plan, Patrizia

      * charge has arrived, busy schedule ahead
      * APS will conduct town meetings Aug/Sep
      * draft Mar/Apr
      * final report October, 2015
      * committee yet to be formed
      * EIC workshop in June at Stonybrook

Readiness Review, Elton

      * Documents:
          1. conduct of operations (COO)
          2. emergency management guidelines (EMG)
          3. experimental safety assessment document (ESAD)
          4. radiation safety assessment document (RSAD)
      * EMG: mainly floor plans, produced, Tom Carstens looking at advanced
        draft
      * ESAD: document sent to Paul Collins and Bert Manzlak, minor
        comments came back from Paul
      * Other documents:
          1. TDR: description of apparatus
          2. instructions for shift takers
           + E. will focus on instructions next
           + Patrizia: can collect OSPs, serve as much of instructions
      * Eugene: summary of equipment checked and updated
           + TDR will be discussed at collaboration meeting, very little in
             repository
           + will re-arrange the structure, streamline it

Calorimetry, Elton

      * FCAL: all detectors working
           + new firmware in FADCs
           + slow controls operations using IU interim system
           + interfaces to slow controls and EPICS need to be done
      * BCAL: understand north side, 1 channel not working
           + looking at one sensor at a time, Mark D. took data
           + south side next
           + then TDCs

Tracking Chambers, Lubomir

      * gas system
           + pure gas, certified, has been delivered
           + ready to start flushing with it
           + Patrizia: ODH OSP not signed yet, now door must be open in gas
             shed if personnel inside
           + will add filters, detectors expensive, gas certified, but a
             guard against mistakes, protects against hydrocarbons and
             water, not oxygen, oxygen not dangerous, will be added later,
             depend on certification alone for now
           + Dave Butler has worked out purging procedure, want to start
             right away

Tracking Chamber Electronics, Beni

      * have latest firmware, can control latency, window size
      * can do one-crate read-out
      * next: cosmic trigger, multiple crates, CDC rate OK, FDC rate 80
        events/hour
      * independent VXS crate coming, then distribute signals to 14 FDC
        crates, will get timing info, plan is for next weeks on this

Tracking Chamber Calibration, Simon

      * ready for CDC, though would like BCAL in trigger
      * can produce efficiencies
      * horizontal cosmics can be analyzed

Trigger, Sascha

      * BCAL
           + global trigger crate
           + now with one crate
           + next 4 crates on south
      * trying playback mode
           + preload, get trigger, works well
           + special trigger type from TS
      * latency 2.4 microseconds, too much
           + 500 ns cable to tagger
           + GTP not in chain
           + on edge of budget
      * next: TDCs
      * TOF
           + 3 crates running
           + can initialize CAEN TDCs, data not read-out yet
      * FDC
           + many crates
           + master TI will not work
           + partition TS no read
           + new trigger crate, 3rd crate in global trigger rack
                o add another DAQ system, need optical fibers still

Data Acquisition, Sergey F.

      * workflow for stand-alone systems working
      * no yet working for common system for several subsystem, problem
        with shared libraries
      * working on debugging and monitoring of DAQ system
           + first version with online histogramming using ROOTSpy

Software Issues, David

      * monitoring problem with connection to ETs started by CODA
      * fixed version of translation tables
           + some wire problems in FDC
      * bench-marking for Chip, ready to run
      * Hall-D/GlueX User Group Meeting talk
      * Open House:
           + walk-through with Keith Welch on platform
           + training May 13, may have dedicated Hall D session in Hall
           + Other displays:
               1. Nick: electronics
               2. Tim: spare straws
               3. Lubomir: cathode plane
           + projector on TOF wall showing event display

Beamline, Alexandre

      * hall in power permit now
      * PSS certification nearly done, after that controlled access
      * beam planned for early afternoon tomorrow
      * dosimetry installed in tagger hall
           + below tagger, upstream of radiator, close to exit wall

Solenoid, Yi

      * power supply, connection issue
           + ran last week up to 800 A to dummy load, ramp rate 2 A/s
           + Joe B. and Mark S. report spikes current fast ramp, contacting
             Danfysik
      * PLC to EPICS running
           + list of control signals for operators being developed
           + list of signals to be archived being developed
      * aim to be ready by June
      * Eugene: Jonathan Creel appointed head of cryo group
      * Cerenkov
           + will get BaBar boxes
           + making list of tasks and resources
                o JLab: readout
                o IU: engineering
                o MIT: deployment to maximize physics
           + need to prepare report to PAC
           + then need to think about funding
           + LDRD proposal submitted, only six this year, ours a bargain

Mark

      * continuing data challenge clean up
           + all JLab ROOT files archived to tape
           + transferring data from UConn for archiving to tape library
             using OSG SRM
           + working with Richard Jones to fix SRM anomalies

Slow Controls, Nerses

      * alarms for MPOD developed
           + same scheme as for CAEN HV, intrinsic differences were worked
             out

Slow Controls, Vanik

      * changes to BCAL voltage control implemented
           + new GUI released
      * tested stage for microscope
           + 3 legs, each leg moves separately
      * EPICS driver for XPS controller
      * TOF HV GUI started

Microscope Fibers, Eugene

      * discussion on possible causes of transmission loss

-- 
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295



More information about the Halld-jlab mailing list